# The Berezin liminf criterion fails for radial Toeplitz operators

### 11 Mar 2026 v1

@@ Abstract
We show that a positive limit inferior of the Berezin transform does not imply essential positivity for radial Toeplitz operators on the Bergman spaces $A^{2}(\mathbb{B}_{d})$ and the Fock spaces $F^{2}(\mathbb{C}^{d})$ in every complex dimension $d\geq 1$. In particular, this disproves the Perälä–Virtanen conjecture in its original one-dimensional Bergman form and shows that the corresponding radial Berezin criterion fails in all dimensions in both the Bergman and Fock settings. For each $d\geq 1$ we construct explicit bounded real-valued radial symbols $f$ for which $\liminf_{|z|\to 1^{-}}\widetilde{f}(z)>0$ on $A^{2}(\mathbb{B}_{d})$ and $\liminf_{|z|\to\infty}\widetilde{f}(z)>0$ on $F^{2}(\mathbb{C}^{d})$, while the essential spectrum of $T_{f}$ contains a negative point. In both settings, the proofs reduce to explicit asymptotics for one-dimensional oscillatory integrals. The examples arise because, for radial symbols, the eigenvalue sequence and the Berezin transform are different asymptotic averages of the same oscillatory symbol, and these averages attenuate the oscillation by different factors.
@@ 1 Introduction
The main result of this paper is that the answer to both Conjecture 1.1 and Question 1.2 is “no” in the radial class in which the conjecture was formulated, not only in dimension one, but actually in all positive dimensions. For every $d\geq 1$ we construct explicit bounded real-valued radial symbols on the Bergman space $A^{2}(\mathbb{B}_{d})$ and the Fock space $F^{2}(\mathbb{C}^{d})$ such that the Berezin transform has strictly positive limit inferior, while the essential spectrum of the corresponding Toeplitz operator contains a negative point. Thus positive Berezin liminf does not characterize essential positivity for radial Toeplitz operators in any complex dimension, on either family of spaces.
@@ 2.2 A radial counterexample on the Fock space
Since $F(0)=\sqrt{\pi}$, by Grönwall’s inequality,

### 16 Mar 2026 v2

# Positive Berezin liminf does not imply essential positivity for radial Toeplitz operators on Bergman and Fock spaces

@@ Abstract
We study whether essential positivity
$\sigma_{\mathrm{ess}}(T_{f})\subset[0,\infty)$
of a radial Toeplitz operator on Bergman and Fock spaces can be detected from the asymptotic behavior of its Berezin transform. For bounded real-valued radial symbols on $A^{2}(\mathbb{D})$, Perälä and Virtanen conjectured that
$\liminf_{|z|\to 1^{-}}\widetilde{f}(z)\geq 0$
should be equivalent to essential positivity, and they asked the analogous question on Fock space. Such a criterion would turn a spectral question into a scalar asymptotic test.
We prove that this fails even in the radial setting in which the conjecture was formulated. For every complex dimension $d\geq 1$, we construct explicit bounded real-valued radial symbols on the Bergman spaces $A^{2}(\mathbb{B}_{d})$ and the Fock spaces $F^{2}(\mathbb{C}^{d})$ whose Berezin transform has strictly positive limit inferior at the boundary (respectively, at infinity), while the essential spectrum of the corresponding Toeplitz operator contains a negative point. In particular, this disproves the Perälä–Virtanen conjecture in its original one-dimensional Bergman form and answers the analogous Fock-space question negatively; more generally, the radial Berezin liminf criterion fails in all dimensions in both settings. The underlying reason is that, for radial symbols, the Toeplitz eigenvalue sequence and the Berezin transform are different asymptotic averages of the same oscillatory symbol, and these averages damp the oscillation by different amounts.
@@ 1 Introduction
The main result of this paper is that the answer to both Conjecture 1.1 and Question 1.2 is “no” in the radial class in which the conjecture was formulated, not only in dimension one, but actually in all positive dimensions. For every $d\geq 1$ we construct explicit bounded real-valued radial symbols on the Bergman space $A^{2}(\mathbb{B}_{d})$ and the Fock space $F^{2}(\mathbb{C}^{d})$ such that the Berezin transform has strictly positive limit inferior, while the essential spectrum of the corresponding Toeplitz operator contains a negative point. Thus positive Berezin liminf does not characterize essential positivity for radial Toeplitz operators in any complex dimension, on either family of spaces, and positivity of $\liminf\widetilde{f}$ is stronger than nonnegativity.
@@ 2.2 A radial counterexample on the Fock space
Since $F(0)=\sqrt{\pi}$,